# A framework for LISA population inference

### v1

@@ A.1 Recovering automatic data selection
$\left[\prod_{i=1}^{N_{\rm obs}}\frac{N_{1}}{K}p_{1}(d_{c,i}|\Lambda)\right]\left(1-\frac{N_{1}}{K}\right)^{K-N_{\rm obs}}\prod_{j=1}^{K-N_{\rm obs}}\left[\frac{N_{2}}{K}p_{2}(d_{c,j}|\Lambda)+\left(1-\frac{N_{2}}{K}\right)p_{n}(d_{c,j})\right].$ (A.20)
We now integrate out the data dependence from the unresolved segments, as we did above. Under the assumptions we have made, the unresolved and noise likelihoods have no support in the resolved source region and so they integrate to unity. The second product is therefore equal to $1$ and the limit of the middle term as $N\rightarrow\infty$ is $\exp(-N_{1})$ as before. So, we are left with
$\left[\prod_{i=1}^{N_{\rm obs}}\frac{N_{1}}{K}\int p_{\rm res}(d_{c,i}|\theta)\alpha(\theta,\Lambda)p(\theta|\Lambda){\rm d}\theta\right]\exp(-N_{1}).$ (A.21)
The rate $N_{1}$ in this model is just $Np_{\rm det}(\Lambda)$ and so this expression looks (almost) correct. The one difference is the presence of the $\alpha(\theta,\Lambda)$ term for each resolved event. We would normally expect the uncorrected likelihood, $p(\mathbf{d}|\Lambda)$ to appear here. However, this is OK. In the model we have had to assume there are two likelihoods for events in the population, and so for given values of the event parameters the likelihood is actually
$p(d_{c}|\theta)=\alpha(\theta,\Lambda)p_{\rm res}(d_{c})|\theta)+(1-\alpha(\theta,\Lambda))p_{\rm b}(d_{c}|\theta).$ (A.22)
For any event that is above threshold, the first term is dominant and this term brings a factor of $\alpha(\theta,\Lambda)$. We therefore conclude that under the contrived assumptions described here we can recover the LVK limit exactly.
The reason that the example is so contrived is that the two formalisms are fundamentally different in philosophy. In the formalism described here, we use all of the data in the inference, and weight segments according to the evidence that they contain a signal. In the standard LVK analysis, a pre-processing is applied to sub-select some of the data. This is done for pragmatic reasons, as the computational cost of doing parameter estimation on all of the LVK data is prohibitive. In the LISA case, we have a signal data set containing many sources at all times and so we are inevitably led to analyse all of the data together. If we did apply the formalism presented here to the LVK case, for example by sampling simultaneously in the parameters of the model, and the flags indicating which likelihood is being used for each segment, then we would find that each segment would appear in the resolved source part of the likelihood a proportion of the time that is determined by the Bayesian evidence for the resolved model component. If we imagine having observed for a very long time, so that all data segments have occurred many times, each sample would remove some segments and add others so that they continue to appear in the relevant proportions. The included segments would contribute the usual factors of $p(d_{c}|\theta)$ to the joint likelihood and unresolved segments will enter the likelihood directly (but with an approximate likelihood model) and so do not need to be explicitly marginalised. We can therefore interpret what we are doing as a generalisation of the LVK-style analysis which replaces the usual strict data-dependent inclusion/exclusion criterion, with one that is probabilistic, and says that events are included in our analysis with a certain data-dependent probability (related to the Bayesian evidence ratio between the resolved and background/noise models).

### v2

@@ Appendix A LVK limit
$\hskip 56.9055pt\equiv{\rm e}^{-N_{1}}p(\{{d}_{c}\}_{\rm rest}|\Lambda)\prod_{j=1}^{K-n_{\rm det}}\left[1+\frac{N}{(K-N_{1})}\frac{\tilde{p}_{1}({d}_{c,i}|\Lambda)}{q({d}_{c,i}|\Lambda)}\right]$ (A.22)
where in the last line we are defining the “noise-likelihood” term $p(\{d_{c}\}_{\rm rest}|\Lambda)$, which appeared as $\int p(d|\Sigma)p(\Sigma|\Lambda)\ {\rm d}\Sigma$ in Eq. (2.8) and we took the $K\rightarrow\infty$ limit for the first term. We now eliminate the final product term by replacing the term in square brackets with its expectation value. This can be justified by the central limit theorem — the product of a large number of independent and identically distributed random variables becomes increasingly concentrated around its expectation value to the power of the number of observations as that number tends to infinity. The expectation value can be computed as
$1+\frac{K}{K-N\xi_{1}(\Lambda)}\int_{d_{c}\in{\rm ndet}}\left(\frac{N}{K}\tilde{p}_{1}(d_{c}|\Lambda)+\left(1-\frac{N_{1}}{K}\right)q(d_{c}|\Lambda)\right)\frac{N}{K-N_{1}}\frac{\tilde{p}_{1}({d}_{c}|\Lambda)}{q({d}_{c}|\Lambda)}\,{\rm d}d_{c}$ (A.23)
The factor outside the integral is the normalisation that comes from the fact that these segments are restricted to be in the non-detected region and so the expectation is taken over the distribution of data within that region (it is one over the integral of the term in brackets over that region). We see that to $O((N/K)^{2})$ the contribution is
$1+\frac{N}{K}\int_{d_{c}\in{\rm ndet}}\tilde{p}_{1}({d}_{c}|\Lambda)\,{\rm d}d_{c}=1+\frac{N}{K}\left(\frac{N_{1}}{N}-\xi_{1}(\Lambda)\right).$ (A.24)
Replacing each square bracketed term in Eq. (A.22) with this expression and then taking the usual limit $K\rightarrow\infty$ we obtain
${\rm e}^{-N_{1}}p(\{{d}_{c}\}_{\rm rest}|\Lambda){\rm e}^{N_{1}-N\xi_{1}(\Lambda)}=p(\{{d}_{c}\}_{\rm rest}|\Lambda){\rm e}^{-N\xi_{1}(\Lambda)}.$ (A.25)
Putting the pieces together, we get
$p(d|\Lambda)=p(\{{d}_{c}\}_{\rm rest}|\Lambda){\rm e}^{-N\xi_{1}(\Lambda)}\prod_{i=1}^{n_{\rm det}}\int p_{1}(d_{c,i}|\Lambda)N_{1}(\theta|\Lambda),$ (A.26)
which takes the same form as Eq. (2.8) once we make, as before, the identification that $\xi_{1}(\Lambda)\approx\xi(\Lambda)$.
In this way we have shown how to get results that look like the LVK result from our formalism in the appropriate limit. One important thing to note is that the “noise likelihood” term, $p(\{{d}_{c}\}_{\rm rest}|\Lambda)$, in the above equations comes only from the unresolved population, see Eq. (A.18), and the detected term comes from the resolved one, while in Eq. (2.8), the whole population is used in both terms (since the split between resolved and unresolved sources was not explicitly done). In the current LVK setup $\alpha(\theta,\Lambda)\sim 1$ and it can be considered constant (though not exactly 1) over the range of parameter space contributing to the likelihood, so that $p_{1,2}(\theta|\Lambda)\simeq p(\theta|\Lambda)$, and we recover Eq. (A.18). However, this is valid only while the background is below the detection threshold. Once a background is measurable, this split between resolved and unresolved sources should be taken into account and the resolvability function should be estimated from data. As the resolvability threshold will be driven by the goodness-of-fit of the assumed background model to the unresolved population, we expect this to occur at a point where there are multiple resolved sources within each segment, since there will have to be multiple unresolved sources contributing to the background to make it appear stochastic. This is the main reason why the current LVK formalism will need to be modified in the future.
This derivation has illustrated where approximations are made in the LVK approach, in particular the fact that we assume that there are a very large number of data segments, and that resolved sources are rare so that no segment can contain more than one. For future detectors these assumptions will become increasingly poor and so the standard LVK analysis will no longer be useable. The formalism described here can be readily applied to that case, however. The below-threshold segments in the above expression could potentially be handled in other ways, for example using a Gaussian approximation and the central limit theorem. The contribution to the Gaussian variance would then come from both resolved and background events and hence depend on the total rate $N$. In any approach that does not do any marginalisation over unresolved segments the whole product over these segments (as in Eq. (A.22)) can be thought of as $p(\{d_{c}\}_{\rm rest}|\Lambda)$, but in that case the $\exp(-N\xi(\Lambda))$ term that appears in Eq. (2.8) should not be there. This term arises because of the marginalisation over the data segments that are not being used for the resolved source analysis. Once we add the background term we are reintroducing that data and so it has not been marginalised out, therefore this term should no longer appear.
Finally, we note that the separation of segments unto those that definitely have resolved events and others that are ambiguous was only done to connect to the LVK limit, and because it reduces computational cost when events are rare. As rates increase, or signals become longer (as in the LISA case), this separation no longer makes sense as most of the data will be being used anyway and segments may contain more than one resolved event. Again, the formalism described here can be directly applied in any general context without problem. If we sample from the full model described here, including the flags (for example using reversible jump to move between the $f_{1}=0$ and $f_{i}=1$ models), then all the segments will include a resolved source a fraction of the time proportional to the Bayes factor. This can be thought of as a variant of the LVK-type analysis in which data segments are chosen probabilistically, rather than with a fixed threshold. Such an approach is in principle able to extract more information from the data, but at the price of significantly higher computational cost since parameter estimation has to be done in every segment that has non-trivial Bayes factor.